OpenStudy (anonymous):

What is the slope of the line which passes through (1, -2) and (-1, 6)?

hartnn (hartnn):

The slope of the line through points (x1,y1) and (x2,y2) is given by : \(\huge m=\frac{y_1-y_2}{x_1-x_2}\) now,just put the values and find m.

OpenStudy (anonymous):

ohh yea lol thanks i remember now how to do it

hartnn (hartnn):

good! so can you tell me what slope you got ? so that i can verify it..

OpenStudy (anonymous):

4

hartnn (hartnn):

+4 or -4 ??

OpenStudy (anonymous):

+4

hartnn (hartnn):

recheck please. the numerator = -2-6 = -8 the denominator = 1-(-1) = 2 so, m= .... ?

OpenStudy (anonymous):

i know what i did its -4 my bad:(

hartnn (hartnn):

yes, m=-4 is correct :)